A woman is one year more than 4times the age of her daughter. In 4 years she will be 3times as old as her daughter will be then. Find their present ages.
Answer by stanbon(75887)

A woman is one year more than 4times the age of her daughter. In 4 years she will be 3times as old as her daughter will be then. Find their present ages.
========
Equations:
w = 4*d
w+4 = 3(d+4)
------
Substitute for "w" and solve for "d":
4d + 4 = 3d+12
d = 8 (daughter's age)
----
Solve for "w":
w = 4d = 32 yrs (the woman's age)
=========================
Cheers,
Stan H.
=========================